1. Answer the short Question:
a. What is the "law of one price"?
b. Write down the absolute form of PPP and interpret it.
c. Write down the relative form of PPP and interpret it.
d. What is a "Spot Exchange Rate " ?
Answer #1

a)

Law of one price:

Law of one price states that in the when conditions of free trade and price flexibility exists, two identical security or asset should sell at the same price regardless of the location. This law helps in eliminating arbitrage opportunities which can arise due to difference is prices of identical products selling at different locations.

b)

S= P ÷ P*

S is exchange rate between two countries

P is price index for a domestic country

P* is price index for a foreign country

This states that exchange rate between two countries will be same to the ratio of price levels of the two countries.

c)

S1 / S0 = (1 + iy) ÷ (1 + ix)

S1 is the spot exchange rate at the end of the time period

S0 is the spot exchange rate at the beginning of the time period

iy is the expected annualized inflation rate of the foreign country y

ix is the expected annualized inflation rate for domestic country x

Relative purchasing power parity relates the changes in expected inflation rates of both the countries to the changes in theri exchange rates. Relative purchasing power parity considers the changes in price levels of both countires and maintains that exchange rates will change to compensate for inflation differentials.


d)

Spot Exchange Rate

Spot exchange rate is the rate at which one currency will trade for another today. It is the price a person needs to pay in one currency to purchase another.

For example, 1 US dollar = 70 indian rupees. It would take 70 INR to purchase one dollar.